OpenStudy (anonymous):

what is the ( at least, at most mean ) in inequality. Anybody explain about inequality!~!

jimthompson5910 (jim_thompson5910):

"at least" means "the smallest you can go (including that number)"

jimthompson5910 (jim_thompson5910):

example it costs at least $10 to get into the party. So the smallest amount you have to pay is $10, but you could pay more

jimthompson5910 (jim_thompson5910):

"at most" is the opposite and it means the most or highest value allowed example: the temperature will be at most 99 degrees today. So the max temp is 99 degrees

OpenStudy (anonymous):

you have to be at least 21 to buy drinks at this bar

OpenStudy (anonymous):

How can I use to the inequality?? like \[>\le \ge <??\]

OpenStudy (anonymous):

that means clearly that if must be 21 or older, i.e. \(\geq 21\)

OpenStudy (anonymous):

Oh what about the at most?

OpenStudy (anonymous):

i have at most $30 in my pocket, \(x\leq \$ 30\)

jimthompson5910 (jim_thompson5910):

"at least" you use \(\Large \ge\) "at most" you use \(\Large \le\)

jimthompson5910 (jim_thompson5910):

x is at least 30 translates to \[\Large x \ge 30\] y is at most 20 translates to \[\Large y \le 20\]

OpenStudy (anonymous):

sorry everybody! I want to know every symbol the inequality..!! so \[a <b, a >b, a \le b, a \ge b \]

jimthompson5910 (jim_thompson5910):

not sure what you're asking

OpenStudy (anonymous):

Like everything about inequality. everything ^!^

OpenStudy (anonymous):

\(a<b\) "a is less than b"

OpenStudy (anonymous):

\(a<b\) \(a\) is to the left of \(b\) on the number line

OpenStudy (anonymous):

\(a\leq b\) a is less than or equal to b a is at most b a is no more than b whatever your favourite english is

OpenStudy (anonymous):

Yes~~ that was my point ~!! thank you so much!! ^&^

OpenStudy (anonymous):

yw

hero (hero):

I'd rather write it this way: \(30 \ge x\) Means that 30 is the most I can have compared to x \(20 \le x\) Means 20 is the least I should have compared to x